# PersonalVideo: High ID-Fidelity Video Customization without Dynamic and Semantic Degradation

### v1

@@ Abstract
The current text-to-video (T2V) generation has made significant progress in synthesizing realistic general videos, but it is still under-explored in identity-specific human video generation with customized ID images. The key challenge lies in maintaining high ID fidelity consistently while preserving the original motion dynamic and semantic following after the identity injection. Current video identity customization methods mainly rely on reconstructing given identity images on text-to-image models, which have a divergent distribution with the T2V model. This process introduces a tuning-inference gap, leading to dynamic and semantic degradation. To tackle this problem, we propose a novel framework, dubbed PersonalVideo, that applies direct supervision on videos synthesized by the T2V model to bridge the gap. Specifically, we introduce a learnable Isolated Identity Adapter to customize the specific identity non-intrusively, which does not comprise the original T2V model’s abilities (e.g., motion dynamic and semantic following). With the non-reconstructive identity loss, we further employ simulated prompt augmentation to reduce overfitting by supervising generated results in more semantic scenarios, gaining good robustness even with only a single reference image available. Extensive experiments demonstrate our method’s superiority in delivering high identity faithfulness while preserving the inherent video generation qualities of the original T2V model, outshining prior approaches. Notably, our PersonalVideo seamlessly integrates with pre-trained SD components, such as ControlNet and style LoRA, requiring no extra tuning overhead.
Figure 1: Results of the proposed PersonalVideo. Given a few or just one reference image of a specific identity, PersonalVideo can generate high ID-fidelity videos aligned with the prompt and seamlessly integrate pre-trained component in the AIGC community. Project page is available at https://personalvideo.github.io/
footnotetext: Project Leader. Support by Alibaba Research Intern Program.
@@ 1 Introduction
Figure 2: Previous T2V customization supervises the tuning process via reconstructing images on T2I models, suffering from a tuning-inference gap. Differently, we aim to directly apply the supervision on generated videos, which aligns with inference and bridges the gap.
@@
To address these challenges, we propose a novel framework, dubbed PersonalVideo, for ID-specific video generation that can achieve high ID fidelity and maintain original motion dynamics and semantic following with only a few images of an identity given. Based on a pre-trained video generation model, our PersonalVideo injects the identity from given images into some learnable modules through an optimization process.
Different from previous supervising this tuning process via reconstructing images on T2I models, our core insight is applying the supervision directly to videos generated by the T2V model thus bridging the tuning-inference gap as shown in Fig. 2. Specifically, inspired by the successful experience of the identity loss [42, 33] in encoder-based T2I customization [13, 15], we introduce the identity loss to the generated videos, which could reward the T2V model generating videos with high ID fidelity. During the tuning time, we first generate videos from random noises and then align face embeddings extracted from generated videos and reference images via the identity encoder [11]. Building on recent advancements in fast sampling methods [39, 43] for T2V models, this process only requires four denoising timesteps with manageable tuning costs and negligible quality loss.
@@
With the non-reconstructive supervision, we further employ simulated prompt augmentation, which is independent of provided reference images and supervises generated results in more semantic scenarios. Unlike the reconstructive training method, they do not correspond to the references nor are they limited by the number of them, effectively mitigating overfitting and demonstrating strong robustness, even when only a single reference image is available.
To achieve the identity injection with preserved motion dynamics and semantic following, it is also essential to design the learnable modules. Through the experiments on T2V diffusion models, we can observe that: (1) In the early stages of the denoising process, the focus is on restoring layout and motion, while the later stages concentrate on recovering detailed object appearance [8, 31]. (2) The spatial self-attention layers in the pre-trained diffusion model play a crucial role in preserving the geometric and shape details for the identity [28, 38], while the spatial cross-attention layers are primarily responsible for preserving semantic information. Based on the observations, we propose an Isolated Identity Adapter which is injected into the spatial self-attention layer in the only last denoising step of the fast sampling diffusion model.
@@
Qualitative and quantitative experiments demonstrate that our PersonalVideo achieves high ID fidelity and effectively preserves the original T2V model’s capabilities. Benefiting from the scalability of T2V models, it also supports any style-specific fine-tuned model and other conditional inputs, such as poses, which offer valuable flexibility for abundant creation in the AIGC community. Our contributions are summarized as follows:
@@
  • We introduce a novel framework, dubbed PersonalVideo, for video personalization with limited images. To bridge the tuning-inference gap, we propose to directly apply the supervision on the generated videos, achieving high ID-fidelity without dynamic and semantic degradation.
@@
  • With the non-reconstructive supervision, we employ simulated prompt augmentation which is independent of the reference images and robustly improves ID-fidelity, even for just a single image.
  • We introduce an Isolated Identity Adapter to inject the identity and effectively mitigate dynamic and semantic degradation. It can also seamlessly combine with pre-trained SD components, without extra tuning effort.
@@ 2 Related Work
Text-to-Video Customization. The T2V customization presents further challenges compared to the T2I customization due to the temporal motion dynamics involved in videos. Currently, only a limited number of works [30, 17, 47, 46] have undertaken early investigations into this area. MagicMe [30] adopts an ID module based on extended Textual Inversion. However, training under T2I reconstruction supervisiondeviates from the T2V inference, leading to inferior ID fidelity and model degradation. DreamVideo [45] customizes the identity given a few images, but the inconvenience lies in the fact that it necessitates additional videos to provide motion patterns. ID-Animator [17] proposes to encode ID-relevant information with a face adapter, which requires thousands of high-quality human videos for fine-tuning, thereby incurring significant costs associated with dataset construction and model training.
Figure 3: Overview of PersonalVideo framework. To bridge the tuning-inference gap, we directly apply ID supervision on generated videos starting from pure noises. During the optimization, we adopt simulated prompt augmentation to supervise generated results in more semantic scenarios. To preserve original motion dynamics and semantic following, we introduce an Isolated Identity Adapter(SIA) to inject the identity into the spatial self-attention layer only for the last denoising steps.
@@ 3 Preliminaries
Text-to-Video Diffusion Models. Text-to-video diffusion models (T2V) [6, 14, 40, 39] are tailored for generating videos by adapting image diffusion models to handle video data. Specifically, the diffusion model $\epsilon_{\theta}$ aims to predict the added noise $\epsilon$ at each timestep $t$ based on text condition $c$, where $t\in\mathcal{U}(0,1)$ is normalized. The training objective can be simplified as a noise-prediction loss:
@@
ID Customization. ID customization for text-to-image models (T2I) enables pre-trained models to generate images that reflect specific identities while adhering to the provided text descriptions. Typically, they optimize a new word embedding for the user-provided ID or fine-tune the generator to enhance ID fidelity. Similarly to the training objective of pre-trained diffusion models, they adopt the reconstruction supervision with L2 noise-prediction loss to bind the user-provided references with the special token to inject the identity to the pre-trained T2I models.
@@ 4 Methodology
Given few images for a specific identity, our goal is to generate high ID-fidelity videos and preserve original motion dynamics and semantic following. To the end, we propose a novel framework, dubbed PersonalVideo, for high ID fidelity video customization as shown in Fig. 3.
4.1 Non-Reconstructive T2V Customization
@@
To bridge the gap, as shown in Fig. 2, we propose a non-reconstructive framework to directly apply supervision on the generated videos for high ID fidelity without dynamic and semantic degradation. Inspired by the successful experience of the identity loss [42, 33] in encoder-based T2I customization [13, 15], we directly apply the identity loss to the generated videos, which could reward the T2V model generating videos with high ID fidelity. Therefore, during the tuning time, we start from pure noise instead of noised references in the previous reconstructive methods. Then we directly supervise the sampled videos to mimic the identity in the generated videos with that in references using ID loss, which closely aligns with human perception and the distribution of the real world.
Specifically, we use pre-trained ID encoder [11] $\phi$ to precisely extract the ID embeddings of the references and the random $i$-th frame of the sample video. Then we minimize the cosine similarity of them to align the identity effectively and directly. Here we crop the faces and adopt image augmentation techniques like color jitter to make it more robust for limited references. Formally, our training objective is:
$\mathcal{L}_{\text{id}}=\mathbb{E}_{c,i}\left[CosSim\left(\phi(I_{id}),\phi(\text{F-T2V}(z_{T},c,i))\right)\right],$ (2)
where $I_{id}$ are the reference images and $c$ are the text prompts with the specific keyword.
@@ 4.2 Isolated Identity Adapter
To achieve the identity injection with preserved original motion dynamics and semantic following, we introduce an Isolated Identity Adapter as shown in Fig. 3, which effectively addresses the dynamic and semantic degradation.
For motion preservation, we conduct exploratory experiments on video diffusion models to investigate the varying focus on motion at different denoising steps. As shown in Fig. 4, we can observe that the motion of the person is formed in the early denoising step. During these steps, the model tends to restore the layout [8] and motion. In contrast, the later steps focus on recovering of the detailed appearance of the objects [31]. Based on the observation, we propose to inject the identity only in the later denoising steps during training and inference time as shown in Fig. 3, to reduce the influence on the motion generation and mitigate the distribution shift caused by static reference images.
On the other hand, we aim to inject identity while preserving the original semantic following capabilities. Regarding this issue, previous study [28] has observed that, the self-attention layers play a crucial role in preserving the geometric and shape details for the identity while the cross-attention layers are primarily responsible for preserving semantic information. Influenced by CustomDiffusion [24], current T2V methods for subject customization choose to finetune the cross-attention layers. However, it tends to undermine the original semantic space, especially when only a single reference image is available, thereby compromising the editability and flexibility of original T2V models. To the end, we propose an Isolated Identity Adapter to inject the identity only into the self-attention layers to preserve the original semantic space. It adopts the residual path of two low-rank matrices including a down block $A^{\text{down}}\in\mathbb{R}^{d\times r}$ and up block $A^{\text{up}}\in\mathbb{R}^{r\times k}$. For the plug-and-play manner, we freeze the pre-trained diffusion model and only insert the adapter into the spatial self-attention layers. Formally, the updated parameter matrices are
@@
4.3 Simulated Prompt Augmentation
Figure 5: Qualitative Comparison with Previous Methods. As observed, both LoRA and MagicMe suffer from inferior ID fidelity. Besides, MagicMe has a degradation of semantic following, e.g., tuning head. In contrast, our PersonalVideo maintains high ID fidelity and preserve the original motion dynamics and semantic following, which significantly surpasses other methods.
Figure 6: Qualitative Results for Single Reference. While LoRA and MagicMe suffer from the inferior ID fidelity and severe degradation of motion dynamics and semantic following, e.g.lift the dumbell, our PersonalVideo achieves robust customization with high ID fidelity and preserved motion dynamics and semantic following.
Figure 7: Qualitative Comparison with Encoder-based Methods. As observed, both PuLID and ID-Animator suffer from inferior ID consistency. Besides, PuLID has a significant dynamic degradation. In contrast, our PersonalVideo maintains high ID fidelity and preserve the original motion dynamics.
@@ 4.3 Simulated Prompt Augmentation
4.4 Training and Inference
Following previous T2V subject customization methods [45], we adopt a two-step training strategy. In the first step, we freeze the video diffusion model and optimize a textual embedding ‘V’ in few steps using Textual Inversion [12] to achieve a coarse identity personalization and serve as an initialization. Typically, this process does not lead to the dynamic and semantic degradation. Then we train the Isolated Identity Adapter with ID loss directly on the generated videos to enhance further identity details, which achieves high ID fidelity and preserves original motion dynamics and semantic following. Building on recent advancements in fast sampling methods for T2V models like AnimateLCM [39] , we could generate a high-quality video in only four steps. To preserve the motion dynamic, we inject the Isolated Identity Adapter in the last step during the training. Besides, inspired by previous work [35], we also use a weight-space loss to mitigate the overfitting and enhance the diversity. Overall, our training loss is:
$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{id}}+\lambda\|\hat{\theta}-\theta\|_{2}^{2},$ (4)
where $\lambda$ is the coefficient, $\theta$ and $\hat{\theta}$ are the initial and optimized weights.
During the inference time, we generate videos with customized textual embedding and our Isolated Identity Adapter. We only introduce our adapter in the final quarter of the denoising steps, which is consistent with the training process. More details are included in Appendix.
@@ 5 Experiments
Figure 8: Compatibility with customized style LoRAs, i.e., RCNZ Cartoon 3D [2], GuoFengRealMix [1] and GuoFeng  [4].
Figure 9: Controllable generation with ControlNet. PersonalVideo can be seamlessly integrated with conditional inputs such as poses to generate controllable identity-specific videos.
@@ 5.1 Experimental Settings
We utilize AnimateDiff as our text-to-video generation model. Unless stated otherwise, we use Stable Diffusion 1.5 with Realistic Vision [3] during the inference. We use ResNet-100 [16] backbone pre-trained on Glint360K [5] dataset as the face encoder. To demonstrate the superiority of our PersonalVideo, we compare it with Magic-Me [30], a recent identity-specific T2V customization method, as well as LoRA [22], where we initialize the keyword with Textual Inversion [12] in the first stage. Additionally, we compare the results with encoder-based methods like PuLID [15] and ID-Animator [17]. We present the details of baselines and training process in Appendix.
@@ 5.2 Qualitative Results
We provide a qualitative comparison between PersonalVideo and baselines. As shown in Fig. 5, both LoRA and MagicMe suffer from inferior ID fidelity, due to the tuning-inference gap. Besides, MagicMe has a severe misalignment of the prompt, e.g., tuning head. In contrast, our PersonalVideo achieves high ID fidelity and preserves the original motion dynamics and semantic following. To demonstrate the robustness of our method, we also compare for only a single image reference as shown in Fig. 6. The results further underscore the superiority of PersonalVideo, with promising robustness to achieve high ID fidelity and preserve motion dynamics and semantic following.
Additionally, we compare the results with recent encoder-based methods for a single reference in Fig. 7. As observed, the direct integration of PuLID on the T2V model suffers from inferior ID consistency and a significant dynamic degradation due to the distribution gap between the image and video training data. Benefiting from a large amount of additional video training data, ID-Animator maintains motion dynamics but still remains poor identity consistency. In contrast, our PersonalVideo maintains high ID fidelity and preserves the original motion dynamics.

### v2

@@ Abstract
The current text-to-video (T2V) generation has made significant progress in synthesizing realistic general videos, but it is still under-explored in identity-specific human video generation with customized ID images. The key challenge lies in maintaining high ID fidelity consistently while preserving the original motion dynamic and semantic following after the identity injection. Current video identity customization methods mainly rely on reconstructing given identity images on text-to-image models, which have a divergent distribution with the T2V model. This process introduces a tuning-inference gap, leading to dynamic and semantic degradation. To tackle this problem, we propose a novel framework, dubbed PersonalVideo, that applies a mixture of reward supervision on synthesized videos instead of the simple reconstruction objective on images. Specifically, we first incorporate identity consistency reward to effectively inject the reference’s identity without the tuning-inference gap. Then we propose a novel semantic consistency reward to align the semantic distribution of the generated videos with the original T2V model, which preserves its dynamic and semantic following capability during the identity injection. With the non-reconstructive reward training, we further employ simulated prompt augmentation to reduce overfitting by supervising generated results in more semantic scenarios, gaining good robustness even with only a single reference image. Extensive experiments demonstrate our method’s superiority in delivering high identity faithfulness while preserving the inherent video generation qualities of the original T2V model, outshining prior methods.
Figure 1: Results of PersonalVideo. Given the reference images of a specific identity, PersonalVideo can generate high ID-fidelity videos with promising motion dynamics and prompt following.
* Project Leader
Figure 2: Analysis of the tuning-inference gap. Previous T2V customization supervises the tuning process via reconstructing images on T2I models, suffering from a tuning-inference gap. Differently, we aim to directly apply the supervision on generated videos, which aligns with inference and bridges the gap.
@@ 1 Introduction
To address these challenges, we propose a novel framework, dubbed PersonalVideo, for ID-specific video generation that can achieve high ID fidelity and maintain original motion dynamics and semantic following with only a few images of an identity given. Different from previous methods with a reconstructing objective on T2I models, our core insight is applying the direct reward to videos generated by the T2V model thus bridging the tuning-inference gap as shown in Fig. 2. Specifically, we integrate feedback from a mixture of differentiable reward models to inject an identity consistent with the reference while preserving the original model’s dynamic and semantic following capability.
Inspired by the successful experience of the identity supervision [39, 32] in encoder-based T2I customization [13, 15], we propose to apply Identity Consistency Reward (ICR) to a randomly selected frame in the generated video, leveraging the similarity feedback from an identity recognition model to consistently inject the reference’s identity without the tuning-inference gap. However, applying only ID reward still suffers from degraded dynamic and semantic following capabilities, which is caused by the inherent distribution shift introduced by the limited number of static images during ID injection. To address this shift, we further propose a Semantic Consistency Reward (SCR) to preserve the original model’s semantic distribution, thereby mitigating the degradation of dynamic and semantic following capabilities. Leveraging the semantic reward model, we can evaluate the image-text correspondence of the sampled video frames and then align the score distribution between the original video and the target video, thus effectively preserving the dynamics and semantics of the original generator without affecting the injection of the identity.
@@
With the non-reconstructive reward supervision, we also employ simulated prompt augmentation, which is independent of provided reference images and supervises generated results in more semantic scenarios. Unlike the reconstructive training method, they do not correspond to each reference nor are they limited by the number of reference images, effectively mitigating overfitting and demonstrating strong robustness, even when only a single reference image is available.
@@
To achieve the identity injection with preserved motion dynamics and semantic following, it is also essential to design the learnable modules. Based on the observations for the denoising steps, we further propose an Isolated Identity Adapter injected in the only later denoising step, which minimizes the impact on the original video’s dynamic. Qualitative and quantitative experiments demonstrate that our PersonalVideo achieves high ID fidelity and effectively preserves the original T2V model’s capabilities. Benefiting from the scalability of T2V models, it also supports any style-specific fine-tuned model, which offer valuable flexibility for abundant creation in the AIGC community. Our contributions are summarized as follows:
@@
  • We introduce a novel framework, dubbed PersonalVideo, for video personalization with static images. To bridge the tuning-inference gap, we propose to apply the direct reward feedback to generated videos, achieving high ID-fidelity without dynamic and semantic degradation.
  • We propose a novel semantic consistency reward to effectively preserve the original model’s semantic distribution without affecting the injection of the identity.
@@
  • With the non-reconstructive supervision, we introduce simulated prompt augmentation which is independent of the reference images and robustly improves ID-fidelity, even for just a single image.
@@ 2 Related Work
Figure 3: Overview of the framework of PersonalVideo. To bridge the tuning-inference gap, we directly apply reward supervision on generated videos starting from pure noises, including identity consistency reward with the reference and semantic consistency reward with the original video. During the optimization, we adopt simulated prompt sampled from the Large Language Model to supervise generated results in more semantic scenarios.
Text-to-Video Customization. The T2V customization presents further challenges compared to the T2I customization due to the temporal motion dynamics involved in videos. Currently, only limited works [29, 17, 43, 42, 47] have undertaken early investigations into this area. MagicMe [29] adopts an ID module based on extended Textual Inversion. However, training under T2I reconstruction supervisiondeviates from the T2V inference, leading to inferior ID fidelity and model degradation. DreamVideo [41] customizes the identity given a few images, but the inconvenience lies in the fact that it necessitates additional videos to provide motion patterns. ID-Animator [17] proposes to encode ID-relevant information with a face adapter, which requires thousands of high-quality human videos for fine-tuning, thereby incurring significant costs associated with dataset construction and model training.
@@ 3 Preliminaries
Text-to-video diffusion models (T2V) [6, 14, 37, 36, 24] are tailored for generating videos by adapting image diffusion models to handle video data. Specifically, the diffusion model $\epsilon_{\theta}$ aims to predict the added noise $\epsilon$ at each timestep $t$ based on text condition $c$, where $t\in\mathcal{U}(0,1)$ is normalized. The training objective can be simplified as a noise-prediction loss:
@@ 4 Methodology
4.1 Non-Reconstructive Reward Framework
@@
To bridge the gap, as shown in Fig. 3, we propose a non-reconstructive framework to directly apply reward on the generated videos for high ID fidelity without dynamic and semantic degradation. Inspired by the successful experience of the identity supervision [39, 32] in encoder-based T2I customization [13, 15], we propose to apply Identity Consistency Reward (ICR) to the generated video. During the tuning time, we start from pure noise instead of noised references in the previous reconstructive methods. Then we directly supervise the a randomly selected frame from the generated video to mimic the identity in the generated videos with that in references using ID similarity reward, which closely aligns with human perception and the distribution of the real world.
Specifically, we use pre-trained ID recognition model [11] $\mathcal{R}_{id}$ to precisely extract the ID embeddings of the references and the random $i$-th frame of the generated video. Then we minimize the cosine similarity of them to align the identity effectively and directly. Here we crop the faces and adopt image augmentation techniques like color jitter to make it more robust for limited references. Formally, our training objective is:
$\mathcal{L}_{\text{ICR}}=\mathbb{E}_{i,c\sim p(c)}\left[\texttt{CosSim}\left(\mathcal{R}_{id}(I_{ref}),\mathcal{R}_{id}(G_{\mathcal{T}}(z_{T},c,i))\right)\right],$ (2)
where $I_{ref}$ are the reference images, $G_{\mathcal{T}}$ is the target trained diffusion model with the VAE decoder, and $c$ is the text prompt with the specific keyword.
4.2 Semantic Consistency Reward
While the non-reconstructive ICR addresses the tuning-inference gap in existing methods, the distribution gap introduced by the limited static references still remains, which makes it challenging to preserve the original dynamics and semantics. To address this issue, we further introduce a novel Semantic Consistency Reward (SCR) to maintain the original model’s semantic distribution, effectively alleviating the decline in dynamic and semantic following. Specifically, we randomly sample a batch of $M$ frames from the video generated by original and target model. Then we leverage the semantic reward model $\mathcal{R}_{sem}$ to evaluate the image-text correspondence of the sampled video frames
$\begin{gathered}V_{c}^{\mathcal{S}}=\texttt{Softmax}(\{\mathcal{R}_{sem}(G_{\mathcal{S}}(z_{T},c,i)\}_{i=1}^{M})),\\ V_{c}^{\mathcal{T}}=\texttt{Softmax}(\{\mathcal{R}_{sem}(G_{\mathcal{T}}(z_{T},c,i)\}_{i=1}^{M})),\end{gathered}$ (3)
where $G_{\mathcal{S}}$ is the frozen original diffusion model with the VAE decoder. Finally, we align the score distribution between the original video and the target video with KL divergence,
$\begin{gathered}\mathcal{L}_{\text{SCR}}=\mathbb{E}_{c\sim p(c)}D_{KL}(V_{c}^{T}||V_{c}^{S}).\end{gathered}$ (4)
Through the distribution alignment with the original generator, we can effectively preserve its dynamics and semantics without affecting the injection of the identity.
Thus, the full learning objective is defined as:
$\begin{gathered}\mathcal{L}_{\text{train}}=\mathcal{L}_{\text{ICR}}+\mathcal{L}_{\text{SCR}}.\end{gathered}$ (5)
@@ 4.4 Isolated Identity Adapter
To achieve the identity injection with preserved motion dynamics and semantic following, it is also essential to design the learnable modules. Based on the observations for the denoising steps shown in Fig. 4, we find that the motion of the person is typically formed in the early denoising step. During these steps, the model tends to restore the layout [8] and motion. In contrast, the later steps focus on recovering of the detailed appearance of the objects [30]. Therefore, we propose to isolatedly inject the identity only in the later denoising steps during training and inference time, to reduce the influence on the motion generation and mitigate the distribution shift caused by static reference images. Formally, it adopts the residual path of two low-rank matrices including a down block $A^{\text{down}}\in\mathbb{R}^{d\times r}$ and up block $A^{\text{up}}\in\mathbb{R}^{r\times k}$. Formally, the updated parameter matrices are
@@
Figure 5: Qualitative comparison for a few references. As observed, both Dreambooth and MagicMe suffer from inferior ID fidelity. In contrast, our PersonalVideo maintains high ID fidelity and preserve the original motion dynamics and semantic following, significantly surpassing others.
@@
Figure 6: Qualitative comparison for a single reference. As observed, the baseline methods not only demonstrate suboptimal identity fidelity but also suffer from compromised dynamic motions and insufficient video quality. In comparison, our PersonalVideo maintains higher ID fidelity without dynamic and semantic degradation, which is consistent with Fig. 5.
@@ 5 Experiments
Figure 7: User study. Our PersonalVideo achieves the best human preference compared with DreamBooth and MagicMe.
Figure 8: Compatibility with customized style LoRAs, i.e., RCNZ Cartoon 3D, GuoFengRealMix, and GuoFeng.
@@ 5.1 Experimental Settings
To demonstrate the effectiveness of our method, we utilize a DiT-based model, HunyuanVideo [24], and a UNet-based model AnimateDiff as our T2V generation models. We use ResNet-100 [16] backbone pre-trained on Glint360K [5] dataset as the identity reward model and HPSv2 [44] as the semantic reward model. To demonstrate the superiority of PersonalVideo, we compare it with MagicMe [29], a recent identity-specific T2V customization method, as well as Dreambooth with LoRA [33, 22]. Additionally, we compare the results with encoder-based methods like IDAnimator [17] and ConsisID [47]. We present the details of baselines and training process in Appendix.
@@ 5.2 Qualitative Results
We provide a qualitative comparison between PersonalVideo and baselines. As shown in Fig. 5, both Dreambooth and MagicMe suffer from inferior ID fidelity, due to the tuning-inference gap. In contrast, our PersonalVideo achieves high ID fidelity and preserves the original motion dynamics and semantic following. To verify the robustness of our method, we also conduct the experiments for only a single image reference and compare with encoder-based methods. As shown in Fig. 6, they not only demonstrate suboptimal identity fidelity but also suffer from compromised dynamic motions and insufficient video quality. The results further underscore the superiority of PersonalVideo, with promising robustness to achieve high ID fidelity and preserve motion dynamics and semantic following. To demonstrate the generalization of our method, we also conduct the experiments on a UNet-based model, Animatediff, in the appendix.